# Thermal avalanches in isolated many-body localized systems

### v2

@@ Abstract
Dynamical analysis.– The impact of thermal region $\mathcal{P}$ on MBL instability should also be observable in the dynamics of the system. We consider the time evolution of EE when the system is initially prepared in the Néel state $\left|\psi(0)\right\rangle{=}\left|\uparrow\downarrow{\cdots}\uparrow\downarrow\right\rangle$. The quantum state at time $t$ is then given by $\left|\psi(t)\right\rangle{=}e^{-itH}\left|\psi(0)\right\rangle$, whose EE is computed by tracing out the spins in half of the system. We use $200{-}400$ disorder realizations (depending on $L$) to get the average EE $\overline{S(t)}$. It is well-known that in the ergodic phase EE grows linearly in time, namley $\overline{S(t)}/S_{\mathrm{P}}{\sim}t$, till it saturates to the volume-law value and, therefore, results in $\overline{S(t)}/S_{\mathrm{P}}{\rightarrow}1$ [86]. On the other hand, in the MBL phase, EE grows logarithmically in time, namely $\overline{S(t)}/S_{\mathrm{P}}{\sim}\ln(t)$ [87], and eventually saturates into an entangled state which may exhibit extensive entanglement, though lower than thermal states [80, 87]. This slow growth of entanglement is due to an extensive set of quasi-local integrals of motions [88]. Interactions produce tiny energy shifts among these quasi-local integrals of motions that dephase product-state superpositions, hence entangling distant regions over exponentially long time [82].
@@
In Fig. 4(a), we plot $\overline{S(t)}/S_{\mathrm{P}}$ as a function of time in a system with disorder strength $W{=}9$ for different choices of $P$. As the results show, for $P{=}0$ where the system is in the MBL phase, EE grows logarithmically in time before saturation to $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}\simeq 0.26$. In contrast, for $P{=}L$, i.e. ergodic phase, EE exhibits linear growth in time which eventually saturates to $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}\simeq 0.74$. To see the impact of the thermal region $\mathcal{P}$, we focus on the saturated value of EE at long time scales. In Fig. 4(b), we plot $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ as a function of $P{/}L$ for $W{=}9$ and various system sizes. Consistent with the static analysis, three distinct regions can be recognized: (i) the early plateau in region I, i.e. $P/L{\lesssim}0.1$, indicates a robust MBL phase which resists against thermalization; (ii) the linear rise of the saturated EE in region II, i.e. $0.60.1{\lesssim}P/L{\lesssim}$ 0.6, determines the intermediate phase where the system tends to thermalize but has not yet reached the volume-law; and (iii) the final region III, i.e. $P/L{\gtrsim}0.6$ where $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ remains almost flat indicating the thermal phase. In region II, the linear increase of $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ becomes sharper in systems of larger sizes suggesting that the intermediate phase vanishes in the thermodynamic limit. These observations are completely in line with our static results.
@@
Conclusion.– The delicate balance between disorder strength and the minimum size of a thermalizing region for triggering the avalanche effect remains a central open question in understanding MBL stability. Here, we systematically address this question in an isolated Heisenberg spin chain subjected to random fields by analyzing EE and GR in both static and dynamical scenarios. Two main findings can be highlighted. First, we show that in isolated systems in order to destabilize the MBL phase the size of the thermal region should increase with system size such that $P/L$ exceeds a threshold value, which our simulations suggest to be $P/L{\gtrsim}0.1$. Second, under strong disorder, enlarging the thermal region drives the system into an intermediate phase before fully thermalizing it. The intermediate phase exhibits hybrid characteristics, neither fully localized nor completely ergodic and can be identified in both static and dynamical analysis in finite systems. The intermediate phase seems to shrink as the system size increases suggesting that in the thermodynamic limit it will vanish.
@@
Acknowledgments.– We acknowledge support from the National Natural Science Foundation of China (Grants No. 12050410253, No. 92065115 and No. 12274059).
Supplementary Materials
I Finite-size scaling analysis
@@
$\overline{\langle O_{L}(W)\rangle}=f(L^{1/\nu}(W-W_{c})),$ (S1)
@@
here, $\nu$ is the universal critical exponent for the second-order phase transition, and $f(\cdot)$ is a dimensionless function whose specific form is unknown. To estimate the optimal values of the critical parameters $W_{c}$ and $\nu$ for our finite-size data, we use the ansatz given by Eq. (S1) and plot our quantity of interest as a function of the scaled disorder strength, i.e. $x{=}L^{1/\nu}(W-W_{c})$, for different system sizes $L$. According to Eq. (S1), these curves collapse to each other if $W_{c}$ and $\nu$ are chosen properly. We use PYTHON package PYFSSA for finite-size scaling analysis (see Refs. [80,81] of the main text) that optimizes an quality function $Q$ as the quality of data collapse and estimate the optimal values of the scaling parameters. The quality function, elaborated in Ref. [82] of the main text, for the quantity $O$ can be written as,
$Q_{O}=\frac{1}{T}\sum_{i,j}\frac{(O_{ij}-\mathcal{O}_{ij})^{2}}{dO_{ij}^{2}+d\mathcal{O}_{ij}^{2}}.$ (S2)
@@
Here, $O_{ij}$ represents the values of the quantity $O$ and $dO_{ij}$ represents the corresponding standard deviations at the scaled disorder strength $x_{ij}{=}L_{i}^{1/\nu}(W_{j}-W_{C})$. Similarly, $\mathcal{O}_{ij}$ and $d\mathcal{O}_{ij}$ respectively represent the estimated values of the master function $f(\cdot)$ and the corresponding standard deviation at $x_{ij}$. The normalization factor $T$ represents the number of terms for which $\mathcal{O}_{ij}$ and $d\mathcal{O}_{ij}$ are defined. The sum in Eq. (S2) includes only those $T$ number of terms.
In the case of an optimal data collapse, the quality function is minimized, i.e., $Q_{O}\sim 1$.
Figure S1: Data collapse of the normalized EE using finite-size scaling analysis for various values of $P$. The values of the critical disorder strength, critical exponent, and the optimal value of the quality function in each case are indicated in the insets of the corresponding plot.
@@
In Figs. S1 and S2, we show the obtained data collapse for the normalized entanglement entropy (EE) and the gap ratio (GR) for various values of $P$, respectively. The optimized values of the critical parameters ($W_{c}$ and $\nu$) and the quality function $Q$ are reported respectively. In the absence of a low-disorder region, for EE, we obtain $W_{c}=4.674$, $\nu=0.646$, and $Q=0.79$. Regarding GR, we obtain $W_{c}=4.882$, $\nu=0.863$, and $Q=0.013$. In Figs. S3(a) and (b), we show the critical disorder strength $W_{c}$ as a function of $P$ in the phase diagram for the EE and the GR, respectively. Both panels show that $W_{c}$ remains almost unchanged for the first few values of $P$ and then increases by enlarging the thermal region $\mathcal{P}$.
Figure S2: Data collapse of the averaged GR using finite-size scaling analysis for various values of $P$. The values of the critical disorder strength, critical exponent, and the optimal value of the quality function in each case are indicated in the insets of the corresponding plot.
Figure S3: (a) averaged EE and (b) averaged GR computed for the system size $L{=}16$ as a function of $W$ and $P$. In both panels, the overlaid dots represent the critical disorder strength $W_{c}$ estimated for $P{=}0,1,{\cdots},9$ using finite-size scaling analysis for system sizes $L{=}10,12,14,\ \text{and}\ 16$.
II phase boundary
To trace out the phase boundaries $W_{c}^{*}(L_{av})$ presented in Fig. 3 of the main text, for a fix $P$, we extract the crossing points of the EE or GR when they are plotted as a function of $W$ for curves of size $L$ and $L+2$. Here $W_{c}^{*}(L_{av})$ is the point where the difference between curves of size $L$ and $L+2$ equals zero. In Fig. S4, we plot $\Delta_{S}=\overline{\langle S(L+2)\rangle}/S_{\mathrm{P}}(L+2)-\overline{\langle S(L)\rangle}/S_{\mathrm{P}}(L)$ as a function of $W$. In each panel, $P$ is fixed. For a given $L$ and $L+2$ we identify $W_{c}^{*}(L_{av})$ as the point where the corresponding $\Delta_{S}$ curve crosses the zero line, indicated by the horizontal dashed line. Similarly, in Fig. S5 we have plotted $\Delta_{r}=\overline{\langle r(L+2)\rangle}-\overline{\langle r(L)\rangle}$ as a function of $W$. The critical disorder strength $W_{c}^{*}(L_{av})$ is estimated using the same method.
To estimate the $(P/L_{av})_{c}^{*}$ phase boundaries, we plot a given quantity as a function of $(P/L)$ for a given disorder strength $W$. The critical point $(P/L_{av})_{c}^{*}$ is estimated as the point where a given quantity curve for a system size $L$ intersects the curve for $L+2$. In the case of EE, $(P/L_{av})_{c}^{*}$ corresponds to the points where $\overline{\langle S(L+2)\rangle}/S_{\mathrm{P}}(L+2)$ = $\overline{\langle S(L)\rangle}/S_{\mathrm{P}}(L)$, and for GR these are the points where $\overline{\langle r(L+2)\rangle}=\overline{\langle r(L)\rangle}$. In Figs. S6 and S7 we have respectively plotted EE and GR as a function of $(P/L)$. In each panel, $W$ is kept fixed as reported. For $W{<}W_{c}^{*}(L_{av})$, the curves for any given quantity and any pair of ($L$, $L{+}2$) do not cross hence, these are not shown.
@@
Figure S4: Difference of the normalized EE for all pairs of $L$ and $L+2$ plotted as a function of the disorder strength for various values of $P$. For a given $L_{av}$, the critical disorder $W_{c}^{*}(L_{av})$ is estimated from the crossing of the corresponding curve with the zero line (indicated by the black dashed line). For a fix $P$, the drifts of $W_{c}^{*}(L_{av})$ from smaller to larger values (corresponding to smaller to larger $L$) is visible.
@@
Figure S5: Difference of the averaged GR for all pairs of $L$ and $L+2$ plotted as a function of the disorder strength for various values of $P$. For a given $L_{av}$, the critical disorder $W_{c}^{*}(L_{av})$ is estimated from the crossing of the corresponding curve with the zero line (indicated by the black dashed line). For a fix $P$, the drifts of $W_{c}^{*}(L_{av})$ from smaller to larger values (corresponding to smaller to larger $L$) is visible.
The results presented in Fig. S6 (Fig. S7) show that for $W{\sim}W_{c}^{*}(L_{av})$ on the MBL side, EE (GR) does not exhibit a plateau and increases linearly with $(P{/}L)$. This indicates that localization is not stable against any small low-disorder region for these values of $W$. A plateau begins to develop as $W$ is increased deep inside the MBL phase, which indicates the robustness of the MBL to a certain size range of the low-disorder region.
@@
Figure S6: Normalized EE as a function of $P/L$ for various values of the disorder strength $W$. From the intersection of two curves corresponding to a pair of $L$ and $L+2$ we estimate $(P/L_{av})_{c}^{*}$ at a given $W$.
@@
Figure S7: Averaged GR as a function of $P/L$ for various values of the disorder strength $W$. From the intersection of two curves corresponding to a pair of $L$ and $L+2$ we estimate $(P/L_{av})_{c}^{*}$ at a given $W$.

### v3

@@ IV Dynamical analysis
The impact of thermal region $\mathcal{P}$ on MBL instability should also be observable in the dynamics of the system. We consider the time evolution of EE when the system is initially prepared in the Néel state $\left|\psi(0)\right\rangle{=}\left|\uparrow\downarrow{\cdots}\uparrow\downarrow\right\rangle$. The quantum state at time $t$ is then given by $\left|\psi(t)\right\rangle{=}e^{-itH}\left|\psi(0)\right\rangle$, whose EE is computed by tracing out the spins in half of the system. In numerical calculations, we employ exact diagonalization to obtain the evolved state and use 200400200{-}400 disorder realizations (depending on $L$) to get average EE $\overline{S(t)}$. It is well-known that in the ergodic phase EE grows linearly in time, namely $\overline{S(t)}/S_{\mathrm{P}}{\sim}t$, till it saturates to the volume-law value and, therefore, results in $\overline{S(t)}/S_{\mathrm{P}}{\rightarrow}1$ [86]. On the other hand, in the MBL phase, EE grows logarithmically in time, namely $\overline{S(t)}/S_{\mathrm{P}}{\sim}\ln(t)$ [26], and eventually saturates into its value for an entangled state which may exhibit extensive entanglement, though lower than thermal states [3, 26].
@@
In Fig. 5(a), we plot $\overline{S(t)}/S_{\mathrm{P}}$ as a function of time in a system with disorder strength W=9W{=}9 for different choices of $P$. As the results show, for P=0P{=}0 where the system is in the MBL phase, EE grows logarithmically in time before saturation to $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}\simeq 0.26$. In contrast, for P=LP{=}L, i.e. ergodic phase, EE exhibits linear growth in time which eventually saturates to $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}\simeq 0.74$. To see the impact of the thermal region $\mathcal{P}$, we focus on the saturated value of EE at long time scales. In Fig. 5(b), we plot $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ as a function of P/LP{/}L for W=9W{=}9 and various system sizes. Consistent with the static analysis, three distinct regions can be recognized: (i) the early plateau in region I, i.e. $P/L{\leq}0.1$, indicates a robust MBL phase which resists against thermalization; (ii) the linear rise of the saturated EE in region II, i.e. $0.1{\leq}P/L{\leq}0.6$, marks the intermediate phase, where the system shows partial thermalization without fully reaching the volume law; and (iii) the final region III, i.e. P/L>0.6P/L{>}0.6 where $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ remains almost flat indicating the thermal phase. In region II, the linear increase of $\overline{S(t{\rightarrow}\infty)}/S_{\mathrm{P}}$ becomes sharper in systems of larger sizes, suggesting that the intermediate phase vanishes in the thermodynamic limit. These observations are completely in line with our static results.
Figure 5: (a) The dynamic of EE in a system of size L=16L{=}16 for different values of $P$, when the system with P=0P{=}0 is determined as localized in the presence of strong disorder W=9W{=}9. (b) the saturated values of the EE, namely $\overline{S(t{\rightarrow\infty)}}{/}S_{P}$, as a function of P/LP{/}L when the size of the system varies and W=9W{=}9. The results are obtained using 200400200{-}400 realizations.
V Conclusion
@@
The delicate balance between disorder strength and the minimum size of a thermalizing region for triggering the avalanche effect remains a central open question in understanding MBL stability. Here, we systematically address this question in an isolated Heisenberg spin chain subjected to random fields by analyzing EE and GR in both static and dynamical scenarios. Two main findings can be highlighted. First, we show that in isolated systems in order to destabilize the MBL phase, the size of the thermal region should increase with system size such that $P/L$ exceeds a threshold value, which our simulations suggest to be P/L>0.1P/L{>}0.1. Second, under strong disorder, enlarging the thermal region drives the system into an intermediate phase before fully thermalizing it. The intermediate phase exhibits hybrid characteristics, neither fully localized nor completely ergodic, and can be identified in both static and dynamical analysis in finite systems. The intermediate phase seems to shrink as the system size increases, suggesting that it vanishes in the thermodynamic limit.
Acknowledgements.
@@
AB acknowledges support from the National Natural Science Foundation of China (grants No. 12274059, No. 12574528 and No. 1251101297).
Appendix A Finite-size scaling analysis
Figure 6: Data collapse of the normalized $E$ using finite-size scaling analysis for various values of $P$. The values of the critical disorder strength, critical exponent, and the optimal value of the quality function in each case are indicated in the insets of the corresponding plot.
Figure 7: Data collapse of the averaged GR using finite-size scaling analysis for various values of $P$. The values of the critical disorder strength, critical exponent, and the optimal value of the quality function in each case are indicated in the insets of the corresponding plot.
Figure 8: (a) averaged EE and (b) averaged GR computed for the system size L=16L{=}16 as a function of $W$ and $P$. In both panels, the overlaid dots represent the critical disorder strength $W_{c}$ estimated for $P{=}0,1,{\cdots},9$ using finite-size scaling analysis for system sizes L=10,12,14,and 16L{=}10,12,14,\ $\text${and}\ 16.
@@ Appendix A Finite-size scaling analysis
$\overline{\langle O_{L}(W)\rangle}=f(L^{1/\nu}(W-W_{c})),$ (2)
@@
here, $\nu$ is the universal critical exponent for the second-order phase transition, and $f(\cdot)$ is a dimensionless function whose specific form is unknown. To estimate the optimal values of the critical parameters $W_{c}$ and $\nu$ for our finite-size data, we use the ansatz given by Eq. (2) and plot our quantity of interest as a function of the scaled disorder strength, i.e. $x{=}L^{1/\nu}(W-W_{c})$, for different system sizes $L$. According to Eq. (2), these curves collapse to each other if $W_{c}$ and $\nu$ are chosen properly. We use PYTHON package PYFSSA for finite-size scaling analysis [83, 84] that optimizes a quality function $Q$ as the quality of data collapse and estimates the optimal values of the scaling parameters. The quality function, elaborated in [85], for the quantity $O$ can be written as,
$Q_{O}=\frac{1}{T}\sum_{i,j}\frac{(O_{ij}-\mathcal{O}_{ij})^{2}}{dO_{ij}^{2}+d\mathcal{O}_{ij}^{2}}.$ (3)
@@
Here, $O_{ij}$ represents the values of the quantity $O$ and $dO_{ij}$ represents the corresponding standard deviations at the scaled disorder strength $x_{ij}{=}L_{i}^{1/\nu}(W_{j}-W_{C})$. Similarly, $\mathcal{O}_{ij}$ and $d\mathcal{O}_{ij}$ respectively represent the estimated values of the master function $f(\cdot)$ and the corresponding standard deviation at $x_{ij}$. The normalization factor $T$ represents the number of terms for which $\mathcal{O}_{ij}$ and $d\mathcal{O}_{ij}$ are defined. The sum in Eq. (3) includes only those $T$ number of terms. In the case of an optimal data collapse, the quality function is minimized, i.e., $Q_{O}\sim 1$.
@@
In Figs. 6 and 7, we show the obtained data collapse for the normalized entanglement entropy (EE) and the gap ratio (GR) for various values of $P$, respectively. The optimized values of the critical parameters ($W_{c}$ and $\nu$) and the quality function $Q$ are reported, respectively. In the absence of a low-disorder region, for EE, we obtain $W_{c}=4.674\pm 0.308$, $\nu=0.646$, and $Q=0.79$. Regarding GR, we obtain $W_{c}=4.871\pm 0.243$, $\nu=0.884$, and $Q=0.013$. In Figs. 8(a) and (b), we show the critical disorder strength $W_{c}$ as a function of $P$ in the phase diagram for the EE and the GR, respectively. Both panels show that $W_{c}$ remains almost unchanged for the first few values of $P$ and then increases by enlarging the thermal region $\mathcal{P}$.
@@
Figure 9: Difference of the normalized EE for all pairs of $L$ and $L+2$ plotted as a function of the disorder strength for various values of $P$. For a given $L_{av}$, the critical disorder $W_{c}^{*}(L_{av})$ is estimated from the crossing of the corresponding curve with the zero line (indicated by the black dashed line). For a fixed $P$, the drifts of $W_{c}^{*}(L_{av})$ from smaller to larger values (corresponding to smaller to larger $L$) is visible.
@@
Figure 10: Difference of the averaged GR for all pairs of $L$ and $L+2$ plotted as a function of the disorder strength for various values of $P$. For a given $L_{av}$, the critical disorder $W_{c}^{*}(L_{av})$ is estimated from the crossing of the corresponding curve with the zero line (indicated by the black dashed line). For a fixed $P$, the drifts of $W_{c}^{*}(L_{av})$ from smaller to larger values (corresponding to smaller to larger $L$) is visible.
@@
Figure 11: Normalized $E$ as a function of $P/L$ for various values of the disorder strength $W$. From the intersection of two curves corresponding to a pair of $L$ and $L+2$ we estimate $(P/L_{av})_{c}^{*}$ at a given $W$.
@@
Figure 12: Averaged GR as a function of $P/L$ for various values of the disorder strength $W$. From the intersection of two curves corresponding to a pair of $L$ and $L+2$ we estimate $(P/L_{av})_{c}^{*}$ at a given $W$.
Appendix B phase boundary
To trace out the phase boundaries $W_{c}^{*}(L_{av})$ presented in Fig. 4 of the main text, for a fixed $P$, we extract the crossing points of the EE or GR when they are plotted as a function of $W$ for curves of size $L$ and $L+2$. Here $W_{c}^{*}(L_{av})$ is the point where the difference between curves of size $L$ and $L+2$ equals zero. In Fig. 9, we plot $\Delta_{S}=\overline{\langle S(L+2)\rangle}/S_{\mathrm{P}}(L+2)-\overline{\langle S(L)\rangle}/S_{\mathrm{P}}(L)$ as a function of $W$. In each panel, $P$ is fixed. For a given $L$ and $L+2$ we identify $W_{c}^{*}(L_{av})$ as the point where the corresponding $\Delta_{S}$ curve crosses the zero line, indicated by the horizontal dashed line. Similarly, in Fig. 10 we have plotted $\Delta_{r}=\overline{\langle r(L+2)\rangle}-\overline{\langle r(L)\rangle}$ as a function of $W$. The critical disorder strength $W_{c}^{*}(L_{av})$ is estimated using the same method.
To estimate the $(P/L_{av})_{c}^{*}$ phase boundaries, we plot a given quantity as a function of $(P/L)$ for a given disorder strength $W$. The critical point $(P/L_{av})_{c}^{*}$ is estimated as the point where a given quantity curve for a system size $L$ intersects the curve for $L+2$. In the case of EE, $(P/L_{av})_{c}^{*}$ corresponds to the points where $\overline{\langle S(L+2)\rangle}/S_{\mathrm{P}}(L+2)$ = $\overline{\langle S(L)\rangle}/S_{\mathrm{P}}(L)$, and for GR these are the points where $\overline{\langle r(L+2)\rangle}=\overline{\langle r(L)\rangle}$. In Figs. 11 and 12 we have respectively plotted EE and GR as a function of $(P/L)$. In each panel, $W$ is kept fixed as reported. For $W{<}W_{c}^{*}(L_{av})$, the curves for any given quantity and any pair of ($L$, L+2L{+}2) do not cross hence, these are not shown.
The results presented in Fig. 11 (Fig. 12) show that for $W{\sim}W_{c}^{*}(L_{av})$ on the MBL side, EE (GR) does not exhibit a plateau and increases linearly with (P/L)(P{/}L). This indicates that localization is not stable against any small low-disorder region for these values of $W$. A plateau begins to develop as $W$ is increased deep inside the MBL phase, which indicates the robustness of the MBL to a certain size range of the low-disorder region.